OpenStudy (anonymous):

Determine the zeros of this function using 2 different methods...

OpenStudy (anonymous):

I used the "expand, foil and factor" method ... what is another way i can determine the zeros of this eq... \[f(x) = (x-5)^{2} -36 \]

OpenStudy (anonymous):

use the identity a^2 -b^2 = (a+b)(a-b) and resolve further

OpenStudy (anonymous):

u get (x+1)(x-11)